# DiCeption: A Generalist Diffusion Model for Visual Perceptual Tasks

### v2

@@ 3.4 Adapting to New Tasks
In practical scenarios, it is often necessary to enable models to quickly adapt to new tasks using only a small amount of training data. Traditional specialized foundation models, however, are typically limited to tasks closely related to their domain and generally require substantial datasets and carefully designed network architectures for adaptation. Diffusion models, while powerful in many respects, cannot easily adapt to downstream tasks through fine-tuning only a few parameters with limited data.
DiCeption effectively addresses this limitation. We conducted experiments on lung segmentation, tumor segmentation, and image highlighting, which represent tasks with varying degrees of overlap with the model’s original domain. We train fewer than 1% of the model’s parameters using LoRA [42] without any complex architectural modifications. Notably, despite the limited availability of training samples (50 per task), DiCeption consistently delivered successful and high-quality performance across all target tasks. These results provide compelling evidence for the potential of DiCeption as a truly unified foundation model.
@@ 4.1 Implementation Details
We randomly select 500k images from the OpenImages [51] dataset and use DepthPro [7] and StableNormal [123] to generate depth and normal data. For point-prompted segmentation, we randomly select 400k images from the SA-1B [49] dataset, as well as 200k images with fine-grained hair masks synthesized from the AM2k [56], AIM500 [57], and P3M-10k [55]. Entity segmentation data is from EntityV2 [81], while semantic segmentation data comes from the COCO-Rem [94], and human pose data is sourced from COCO [62]. For few-shot fine-tuning, we select 50 samples from the publicly labeled Chest X-Ray dataset [108], LOL-v2 [121], and Kaggle’s Brain Tumor dataset as training samples. More details can be found in Appendix  S1.
@@
Inference.
We follow the settings of the pretrained model and perform 28 steps of denoising during inference. The inference can be run on a GPU of 24GB memory with a batch size of 4.
4.2 Comparisons with Existing Methods
Table 1: Quantitative comparison of depth estimation with both specialized models and multi-task models on zero-shot datasets. Our visual generalist model can perform on par with state-of-the-art models. We use the same evaluation protocal ($\dagger$) as Genpercept [114].
Method Training KITTI [32] NYUv2 [74] ScanNet [23] DIODE [102] ETH3D [92]
@@ 4.2 Comparisons with Existing Methods
MiDaS [86] 2M 0.236 0.630 0.111 0.885 0.121 0.846 0.332 0.715 0.184 0.752
Omnidata [27] 12.2M 0.149 0.835 0.074 0.945 0.075 0.936 0.339 0.742 0.166 0.778
DPT-large [85] 1.4M 0.100 0.901 0.098 0.903 0.082 0.934 0.182 0.758 0.078 0.946
DepthAnything [118] 63.5M 0.080 0.946 0.043 0.980 0.043 0.981 0.261 0.759 0.058 0.984
DepthAnything v2 [119] 62.6M 0.080 0.943 0.043 0.979 0.042 0.979 0.321 0.758 0.066 0.983
Depth Pro [7] - 0.055 0.974 0.042 0.977 0.041 0.978 0.217 0.764 0.043 0.974
Metric3D v2 [43] 16M 0.052 0.979 0.039 0.979 0.023 0.989 0.147 0.892 0.040 0.983
DiverseDepth [125] 320K 0.190 0.704 0.117 0.875 0.109 0.882 0.376 0.631 0.228 0.694
LeReS [126] 354K 0.149 0.784 0.090 0.916 0.091 0.917 0.271 0.766 0.171 0.777
HDN [128] 300K 0.115 0.867 0.069 0.948 0.080 0.939 0.246 0.780 0.121 0.833
GeoWizard [31] 280K 0.097 0.921 0.052 0.966 0.061 0.953 0.297 0.792 0.064 0.961
DepthFM [33] 63K 0.083 0.934 0.065 0.956 - - 0.225 0.800 - -
Marigold [47] 74K 0.099 0.916 0.055 0.964 0.064 0.951 0.308 0.773 0.065 0.960
DMP Official [54] - 0.240 0.622 0.109 0.891 0.146 0.814 0.361 0.706 0.128 0.857
GeoWizard [31] 280K 0.129 0.851 0.059 0.959 0.066 0.953 0.328 0.753 0.077 0.940
DepthFM [33] 63K 0.174 0.718 0.082 0.932 0.095 0.903 0.334 0.729 0.101 0.902
Genpercept [114] 90K 0.094 0.923 0.091 0.932 0.056 0.965 0.302 0.767 0.066 0.957
Painter [109] 24K 0.324 0.393 0.046 0.979 0.083 0.927 0.342 0.534 0.203 0.644
Unified-IO [69] 48K 0.188 0.699 0.059 0.970 0.063 0.965 0.369 0.906 0.103 0.906
4M-XL [72] 759M 0.105 0.896 0.068 0.951 0.065 0.955 0.331 0.734 0.070 0.953
OneDiffusion  [53] 500K 0.101 0.908 0.087 0.924 0.094 0.906 0.399 0.661 0.072 0.949
Ours-single 500K 0.081 0.942 0.068 0.949 0.078 0.945 0.267 0.709 0.059 0.969
Ours 500K 0.075 0.945 0.072 0.939 0.075 0.938 0.243 0.741 0.053 0.967
Table 2: Quantitative comparison of surface normal estimation with both specialized models and multi-task models. All methods are evaluated with the same evaluation protocol of StableNormal [123].
Method Training NYUv2 [74] ScanNet [23] DIODE-indoor [102]
Samples $\downarrow$ $\downarrow$ $11.25^{\circ}$$\uparrow$ $22.5^{\circ}$$\uparrow$ $30^{\circ}$$\uparrow$ $\downarrow$ $\downarrow$ $11.25^{\circ}$$\uparrow$ $22.5^{\circ}$$\uparrow$ $30^{\circ}$$\uparrow$ $\downarrow$ $\downarrow$ $11.25^{\circ}$$\uparrow$ $22.5^{\circ}$$\uparrow$ $30^{\circ}$$\uparrow$
DINSE [3] 160K 18.572 10.845 54.732 74.146 80.256 18.610 9.885 56.132 76.944 82.606 18.453 13.871 36.274 77.527 86.976
Geowizard [31] 280K 20.363 11.898 46.954 73.787 80.804 19.748 9.702 58.427 77.616 81.575 19.371 15.408 30.551 75.426 86.357
GenPercept [114] 90K 20.896 11.516 50.712 73.037 79.216 18.600 8.293 64.697 79.329 82.978 18.348 13.367 39.178 79.819 88.551
Marigold [47] 90K 20.864 11.134 50.457 73.003 79.332 18.463 8.442 64.727 79.559 83.199 16.671 12.084 45.776 82.076 89.879
StableNormal [123] 250K 19.707 10.527 53.042 75.889 81.723 17.248 8.057 66.655 81.134 84.632 13.701 9.460 63.447 86.309 92.107
Unified-IO [68] 210K 28.547 14.637 39.907 63.912 71.240 17.955 10.269 54.120 77.617 83.728 31.576 16.615 27.855 64.973 73.445
4M-XL [72] 759M 37.278 13.661 44.660 60.553 65.327 30.700 11.614 48.743 68.867 73.623 18.189 12.979 36.622 81.844 87.050
Ours-single 500K 18.267 10.238 52.393 76.802 83.113 19.892 12.424 45.930 74.341 81.965 17.611 8.912 62.030 80.827 86.474
Ours 500K 18.302 10.538 52.533 75.977 82.573 19.348 12.129 46.410 74.805 82.176 17.946 8.686 62.641 81.152 85.398
@@
Method AR-$\uparrow$ AR-$\uparrow$ AR-$\uparrow$
EntityV2 [81] 0.313 0.551 0.683
@@
Table 4: Evaluation of human keypoints estimation on MS COCO.
HRNet[96] HRFormer[129] ViTPose[115] Painter[109] Ours
[…]
Table 5: Evaluation of semantic segmentation on the MS COCO (with category ID).
Method SparK [98] OneFormer [44] Mask2Former [18] Ours
[…]
Figure 2: Comparisons of mIoU with SAM-vit-h. We achieve results on par with SAM using only 0.06% of their data (600K vs.  1B).
We compare the performance of specialized models, existing multi-task models, and our DiCeption across various tasks. Specifically, we evaluate depth using the same protocol as Genpercept, normal estimation using the same method as StableNormal, point-prompted segmentation using the same approach as SAM, and human keypoints using the same method as Painter. We also assess semantic and entity segmentation on the MS COCO dataset. For entity segmentation, we assigned all predicted categories to the same label.
As in Tables 1 and 2, our DiCeption significantly outperforms existing multi-task models and achieves performance on par with state-of-the-art specialized models or demonstrates only an acceptable performance decrease. For point-prompted segmentation, as shown in Figure 2, we achieve results on par with SAM-vit-h using only 0.06% of their data. SAM shows a clear advantage only on certain out-of-distribution datasets that are outside the scope of our model’s training, such as WoodScape fisheye dataset. Notably, while most specialized models require extensive data or complex data pipelines, our method achieves excellent results with significantly less data, the majority of which is obtained through simple random selection. For entity segmentation, we observe that our model performs poorly on small objects. We believe that this is due to the limited amount of data, which causes the model to overfit larger objects. This problem can be solved by introducing more data towards small objects.
We observe that, although our model generates high-quality visualizations for human pose and semantic segmentation, the corresponding evaluation metrics remain relatively low. For human keypoints, this is primarily due to two factors: Firstly, we utilize skeletal-form RGB images rather than heatmaps. While the former produces visually appealing results, the extraction of keypoints during post-processing introduces considerable errors. Secondly, our evaluation follows the 192$\times$256 top-down human keypoints protocol. The original 192$\times$256 images are resized to 768$\times$768 before being input into the model, resulting in extremely blurred inputs that likely contribute to the diminished performance.
Semantic segmentation also introduces considerable errors during post-processing. We offer a comprehensive explanation of the metric degradation in the Appendix B.3. For individual images, we can adjust hyperparameters during the post-processing of RGB mask to achieve the optimal masks. However, different results require different optimal hyperparameters, and manually adjusting each sample in the validation set during evaluation is impractical due to the excessive labor involved. This leads to remarkable semantic segmentation visualizations produced by our method but relatively low metrics. We achieve excellent results by training solely on COCO-Rem. Furthermore, our semantic segmentation can distinguish different instances of the same semantic category. We believe we have strongly demonstrated the potential of our method for instance-level semantic segmentation with IDs. Please check Appendix B for more details about post-processing.
4.3 Comparisons with Our Single-task Models
For the training of single-task models, we ensure that the network architecture remains the same and the total amount of training data seen for each specific task is the same as that for the multi-task model. For example, if the multi-task model trains for 100 iterations with 4 depth data samples per batch, the single-task model will also be trained for 100 iterations with 4 data samples per batch. In our current data setting (approximately 1.8 million samples), we have not observed a significant gap between the multi-task and single-task models, nor have we seen a trend of mutual promotion between different tasks, as shown by “Ours-single” in Tables 1, 2, 3 and Figure 2.
4.4 Ablations
Multi-point Prompted Segmentation.
Ambiguity is a significant issue in point-prompted segmentation. For example, if a point is placed on a person’s clothing, the model may segment the clothing, but the desired result is the person. Therefore, more points are needed to resolve this ambiguity. As illustrated in Figure 3, additional points help the model better segment the desired results.
Figure 3: Comparisons between 1-point and 5-point segmentation of mIoU on all 23 validation datasets.
Architecture of Diffusion Model.
Before the advent of DiT [77], the UNet architecture was predominantly used in diffusion models. We also conduct multi-task experiments based on a UNet pretrained model SDXL [78]. Specifically, we follow Marigold’s [47] approach by expanding the first convolution layer’s input channels from 4 to 8 to accommodate image inputs, and similarly use task prompts to guide the model in solving different tasks. However, as shown in Figure S16, we find that this approach is not sufficiently effective, even for a minimal multi-task scenario involving only depth and normal estimation. We attribute this limitation to the superior information aggregation capabilities of the Transformer architecture compared to UNet. While the UNet suffers from significant information loss during downsampling, the Transformer maintains more comprehensive representations, enabling it to perform better in multi-task scenarios.

### v3

@@ 3.4 Adapting to New Tasks
Practical applications often require models to adapt quickly to new tasks with limited training data. Traditional foundation models, however, are often domain-specific and require extensive data and architectural modifications for adaptation. Powerful diffusion models also struggle with efficient adaptation to downstream tasks via few-parameter fine-tuning on limited data.
DiCeption effectively addresses this limitation. We conducted experiments on lung segmentation, tumor segmentation, and image highlighting, which represent tasks with varying degrees of overlap with the model’s original domain. We train fewer than 1% of the model’s parameters using LoRA hu2021lora without any complex architectural modifications. Notably, despite the limited availability of training samples (50 per task), DiCeption consistently delivered successful and high-quality performance across all target tasks. These results provide compelling evidence for the potential of DiCeption as a unified foundation model.
@@ 4.1 Implementation Details
We randomly select 500k images from the OpenImages kuznetsova2020open dataset and use DepthPro bochkovskii2024depth and StableNormal ye2024stablenormal to generate depth and normal annotations. For interactive segmentation, we randomly select 400k images from the SA-1B kirillov2023segment dataset, as well as 200k images with fine-grained hair masks synthesized from the AM2k li2022bridging , AIM500 li2021deep , and P3M-10k li2021privacy . Entity segmentation data is from EntityV2 qi2022high , while instance segmentation data comes from the COCO-Rem singh2024benchmarkingobjectdetectorscoco , and human pose data is sourced from COCO lin2015microsoftcococommonobjects . For few-shot fine-tuning, we select 50 samples from the Chest X-Ray dataset wang2017hospital , LOL-v2 yang2020fidelity , and Kaggle’s Brain Tumor dataset as training samples. More details can be found in Appendix  A.
@@
Our training lasts for 24 days using 4 NVIDIA H800 GPUs. We employ the AdamW optimizer with a constant learning rate of 2ee5-5 and a batch size of 28 per GPU. We found that the training process is highly stable. However, the convergence speed for segmentation tasks was slower compared to depth and normal tasks. Therefore, we increased the proportion of segmentation data in each batch. Specifically, in each batch, depth and normal each account for 15%, interactive segmentation, entity segmentation, and instance segmentation each account for 20%, and pose estimation accounts for 20%. We observe that, by the end of training, despite the loss no longer significantly decreasing, the model’s performance on segmentation tasks continues to improve.
During few-shot fine-tuning, we apply a rank-128 LoRA to all attention $Q$, $K$, and $V$ layers in the network, which accounts for less than 1% of the total network parameters. The task prompts for different tasks are “image-to-segmentation lung," “image-to-segmentation tumor," and “image-to-highlight." LoRA training is conducted on a single NVIDIA H100 GPU, with a constant learning rate of 2ee5-5 and a batch size of 8. Please refer to Appendix  D for more few-shot fine-tuning visualizations.
Inference.
We perform 28 steps of denoising during inference which follows the settings of the pre-trained model SD3 esser2024scaling . The inference can be run on a GPU of 24GB memory with a batch size of 4. The classifier-free-guidance value is by default set to 2, more analysis in Appendix B.
4.2 Comparisons with Existing Methods
Table 1: Quantitative comparison of depth estimation with both specialized models and multi-task models on zero-shot datasets. Our visual generalist model can perform on par with SOTA models. We use the same evaluation protocol ($\dagger$) as Genpercept xu2024diffusion .
Method Training KITTI kitti NYUv2 nyu ScanNet scannet DIODE diode ETH3D eth3d
@@ 4.2 Comparisons with Existing Methods
MiDaS midas 2M 0.236 0.630 0.111 0.885 0.121 0.846 0.332 0.715 0.184 0.752
Omnidata omnidata 12.2M 0.149 0.835 0.074 0.945 0.075 0.936 0.339 0.742 0.166 0.778
DPT-large dptlarge 1.4M 0.100 0.901 0.098 0.903 0.082 0.934 0.182 0.758 0.078 0.946
DepthAnything yang2024depth 63.5M 0.080 0.946 0.043 0.980 0.043 0.981 0.261 0.759 0.058 0.984
DepthAnything v2 yang2024depth2 62.6M 0.080 0.943 0.043 0.979 0.042 0.979 0.321 0.758 0.066 0.983
Depth Pro bochkovskii2024depth - 0.055 0.974 0.042 0.977 0.041 0.978 0.217 0.764 0.043 0.974
Metric3D v2 hu2024metric3d 16M 0.052 0.979 0.039 0.979 0.023 0.989 0.147 0.892 0.040 0.983
DiverseDepth diversedepth 320K 0.190 0.704 0.117 0.875 0.109 0.882 0.376 0.631 0.228 0.694
LeReS leres 354K 0.149 0.784 0.090 0.916 0.091 0.917 0.271 0.766 0.171 0.777
HDN hdn 300K 0.115 0.867 0.069 0.948 0.080 0.939 0.246 0.780 0.121 0.833
GeoWizard fu2024geowizard 280K 0.097 0.921 0.052 0.966 0.061 0.953 0.297 0.792 0.064 0.961
DepthFM gui2024depthfm 63K 0.083 0.934 0.065 0.956 - - 0.225 0.800 - -
Marigold ke2024repurposing 74K 0.099 0.916 0.055 0.964 0.064 0.951 0.308 0.773 0.065 0.960
DMP Official lee2024exploiting - 0.240 0.622 0.109 0.891 0.146 0.814 0.361 0.706 0.128 0.857
GeoWizard fu2024geowizard 280K 0.129 0.851 0.059 0.959 0.066 0.953 0.328 0.753 0.077 0.940
DepthFM gui2024depthfm 63K 0.174 0.718 0.082 0.932 0.095 0.903 0.334 0.729 0.101 0.902
Genpercept xu2024diffusion 90K 0.094 0.923 0.091 0.932 0.056 0.965 0.302 0.767 0.066 0.957
Painter wang2023images 24K 0.324 0.393 0.046 0.979 0.083 0.927 0.342 0.534 0.203 0.644
Unified-IO lu2022unified 48K 0.188 0.699 0.059 0.970 0.063 0.965 0.369 0.708 0.103 0.906
4M-XL mizrahi20234m 759M 0.105 0.896 0.068 0.951 0.065 0.955 0.331 0.734 0.070 0.953
OneDiffusion  le2024diffusiongenerate 500K 0.101 0.908 0.087 0.924 0.094 0.906 0.399 0.661 0.072 0.949
Ours-single 500K 0.064 0.952 0.066 0.953 0.077 0.942 0.283 0.717 0.052 0.971
Ours 500K 0.069 0.949 0.061 0.960 0.072 0.944 0.289 0.722 0.050 0.975
Table 2: Quantitative comparison of surface normal estimation with both specialized models and multi-task models. All methods are evaluated with the same method of StableNormal ye2024stablenormal .
Method Training NYUv2 nyu ScanNet scannet DIODE-indoor diode
@@
DINSE dinse 160K 18.572 10.845 54.732 74.146 80.256 18.610 9.885 56.132 76.944 82.606 18.453 13.871 36.274 77.527 86.976
Geowizard fu2024geowizard 280K 20.363 11.898 46.954 73.787 80.804 19.748 9.702 58.427 77.616 81.575 19.371 15.408 30.551 75.426 86.357
GenPercept xu2024diffusion 90K 20.896 11.516 50.712 73.037 79.216 18.600 8.293 64.697 79.329 82.978 18.348 13.367 39.178 79.819 88.551
Marigold ke2024repurposing 90K 20.864 11.134 50.457 73.003 79.332 18.463 8.442 64.727 79.559 83.199 16.671 12.084 45.776 82.076 89.879
StableNormal ye2024stablenormal 250K 19.707 10.527 53.042 75.889 81.723 17.248 8.057 66.655 81.134 84.632 13.701 9.460 63.447 86.309 92.107
Unified-IO lu2024unified 210K 28.547 14.637 39.907 63.912 71.240 17.955 10.269 54.120 77.617 83.728 31.576 16.615 27.855 64.973 73.445
4M-XL mizrahi20234m 759M 37.278 13.661 44.660 60.553 65.327 30.700 11.614 48.743 68.867 73.623 18.189 12.979 36.622 81.844 87.050
Ours-single 500K 18.292 10.145 52.693 76.966 83.041 18.807 10.327 52.919 75.152 82.968 16.229 11.012 50.137 83.573 88.972
Ours 500K 18.338 10.106 52.850 77.079 82.903 18.842 10.266 53.610 74.895 82.864 16.297 11.117 50.548 83.325 88.774
Table 3: Evaluation of human keypoints estimation on MS COCO.
HRNethrnet HRFormerhrformer ViTPosevitpose Painterwang2023images Ours
@@
Figure 3: Comparisons of mIoU with SAM-vit-h. We achieve results on par with SAM using only 0.06% of their data (600K vs. 1B). The performance of SAM is clearly better only on some datasets that are out-of-distribution for us, such as the Woodscape woodscape Fisheye dataset.
We compare the performance of specialized models, existing multi-task models, and our DiCeption across various tasks. Specifically, we evaluate depth using the same protocol as Genpercept xu2024diffusion , normal estimation using the same method as StableNormal ye2024stablenormal , interactive segmentation using the same approach as SAM ravi2024sam , and human keypoints using the same method as Painter wang2023images . We also assess instance segmentation and entity segmentation on the MS COCO dataset. For entity segmentation, we assigned all predicted categories to the same label.
Table 4: Evaluation of text-based instance segmentation on the MS COCO.
Method SparK spark OneFormer oneformer Mask2Former mask2former Ours
[…]
As in Tables 1 and 2, our DiCeption outperforms existing multi-task models and achieves performance on par with state-of-the-art specialized models or demonstrates only an acceptable performance decrease. Although some multi-task methods achieve marginally better performance on certain datasets, such as Painter wang2023images and Unified-IO lu2024unified , they exhibit considerably poorer results on others such as outdoor settings (KITTI) and NYUv2 normal map benchmark. This further underscores the robust generalization capabilities of our approach. We contend that focusing on a model’s performance across diverse datasets is more meaningful, as it better reflects the model’s generalization ability and real-world applicability.
For interactive segmentation, as shown in Figure 3, we achieve results on par with SAM-vit-h using only 0.06% of their data. SAM shows a clear advantage only on certain out-of-distribution datasets that are outside the scope of our model’s training, such as WoodScape fisheye dataset. Notably, while most specialized models require extensive data or complex data pipelines, our method achieves excellent results with significantly less data and no training data cherry-picking. Evaluation across diverse datasets highlights the strong in-the-wild generalization capability of our model, demonstrating that it does not overfit to the biases inherent in specific datasets.
We observe that, although our model generates high-quality visualizations for human pose and instance segmentation, the corresponding evaluation metrics remain relatively low. This is also observed on the evaluation of small objects in entity segmentation. We found that this is due to the errors introduced by the post-processing rather than our model’s performance. In Appendix  C, we provide a comprehensive explanation of the post-processing procedure and analyze the underlying causes of metrics degradation.
4.3 Ablations and Analysis
Model designs, classifier-free guidance and pixel-aligned training.
Our crucial analyses covering the elucidation of critical designs for effectively re-purposing diffusion models for perception tasks, as well as significant findings and insights, are detailed in the Appendix due to space limit. Specifically, the analysis of different architectures and input paradigms is presented in Appendix B.1B.2 and  B.3. The effectiveness of modest classifier-free guidance in improving results is discussed in Appendix B.4. The inherent few-step capability of flow-matching on perception tasks is analyzed in Appendix B.5. The benefits of pixel-aligned training are detailed in Appendix B.6 and B.7.
Comparisons with Our Single-task Models.
For the training of single-task models, we ensure that the network architecture remains the same and the total amount of training data seen for each specific task is the same as that for the multi-task model. For example, if the multi-task model is trained for 100 iterations with 4 depth data samples per batch, the single-task model will also be trained for 100 iterations with 4 data samples per batch. In our current data setting (approximately 1.8 million samples), we have not observed a significant gap between the multi-task and single-task models, nor have we seen a trend of mutual promotion between different tasks, as shown by “Ours-single" in Tables 1, 2, 6 and Figure 3.
Multi-point Prompted Segmentation.
Ambiguity is a significant issue in interactive segmentation. For example, if a point is placed on a person’s clothing, the model may segment the clothing, but the desired result is the person. Therefore, more points are needed to resolve this ambiguity. As illustrated in Table 5, additional points help the model better segment the desired results.
Table 5: Comparisons between 1-point and 5-point as input. 5 points are selected randomly.
Method 1-point 5-point
[…]
@@ 4.3 Ablations and Analysis
[…]
EntityV2 qi2022high 0.313 0.551 0.683
@@
One-step Training and One-step Inference.
Genpercept xu2024diffusion demonstrates that diffusion model trained with one-step denoising significantly enhances both the speed and accuracy of perceptual tasks. However, our experimental results reveal a notable increase of failure cases when applying one-step diffusion in a multi-task setting, as illustrated in Figure 4. We believe that this is due to the potential overlap of denoising trajectories for different tasks. These overlapping trajectories can interfere with each other, resulting in failure cases with one-step inference. In contrast, in a single-task setting, since the denoising trajectories pertain to a single task, one-step is more effective and stable. However, we observe that our model, trained with multi-step denoising, can be applied directly to few-step inference with minimal degradation in performance. We provide the results and a more detailed analysis in Appendix B.5 due to space limitations.
Figure 4: The model trained with 1-step denoising tends to produce more failure cases in multi-task scenarios.